OpenStudy (anonymous):

Square GHIJ is inscribed in circle K. If GH = 46, calculate the area of the three shaded segments formed by the circle and square.

OpenStudy (anonymous):

i can do this

OpenStudy (anonymous):

lalaly i need u!!!haha.....someone please help

OpenStudy (anonymous):

wouldnt it be 4 regions left when u put a square

OpenStudy (anonymous):

no i gotta go - see you guys tomorrow

OpenStudy (dumbcow):

well the radius if 23sqrt(2) because it forms half the diagonal of the square Area of square = 46^2 Area of circle = pi*(23sqrt(2))^2 = 2*pi*23^2

OpenStudy (anonymous):

radius is 23sqrt(2)

OpenStudy (anonymous):

1058 pie -48^2

OpenStudy (anonymous):

1019.81

OpenStudy (anonymous):

It has 3 of 4 sides shaded and im trying to calculate the area of the 3 that are shaded.

OpenStudy (anonymous):

764.86

OpenStudy (lalaly):

JH is diameter JH=HI = 46 ... by pythagorus so JH^2 = 46^2 + 46^2 JH = 65.053 Radius = diameter/2 = 32.526 Area of circle = pi*r^2 = 3.14*1058=3322.12 subtractin area of square frm area of circle Area of 4Parts = 1206.12 area of 1 Part = 1206/4 = 301.15 area of 3 Parts = 301.15*3 = 904.59

OpenStudy (dumbcow):

ok then take (AreaCircle - AreaSquare)*(3/4)

OpenStudy (anonymous):

lalay got it!!!!.....you are soo smart...thx for helping everybody....i still have a few more;)

OpenStudy (dumbcow):

i get 905.85, lala you could be off due to rounding